OpenStudy (anonymous):

What is the slope of the line through the points (2, 0) and (3, -3)? m = _?_

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( -3- 0)/(3-2) m = ???

OpenStudy (anonymous):

is -3 y1?

jimthompson5910 (jim_thompson5910):

(2,0) is (x1,y1) so y1 = 0

OpenStudy (anonymous):

ok finally fifnished... got 0

OpenStudy (anonymous):

as my final answer?!

OpenStudy (anonymous):

Your slope is -3

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( -3- 0)/(3-2) m = (-3)/1 m = -3 So the slope is -3

OpenStudy (anonymous):

thank you, thank you thank you